OpenStudy (anonymous):

A field is to be fertilized at a cost of $0.08 per square yard. The rectangular part of the field is 95 yards long and the diameter of each semicircle is 49 yards. Find the cost of fertilizing the field. Use 3.14 for \[\Pi\]

OpenStudy (anonymous):

u hv not given the figure but.......... the rectangular field has dimensions 95 yrd and 49 yrd so area of rectangular part = L x B = 95 yrd x 49 yrd = 4655 sqyrd diameter of semicircle is 49 yrd => radius = dia/2 = 49/2 yrd = 24.5 yrd area of full circle = π r^2 π r^2 so area of semi-circle = ------- 2 π r^2 so area of two semi-circles is = 2 * ------- = πr^2 = 3.14 * (24.5)^2 = 3.14 * 600.25 2 so area of two semi-circles is = 1884.79 sq yrd

OpenStudy (anonymous):

this gives total area of the field = 4655 sqyrd + 1884.79 sqyrd = 6539.79 sqyrd cost of fertilising 1 sqyrd = $ 0.08 so cost of fertilising 6539.79 sqyrd = 6539.79 * $ 0.08 = $ 523.18